# Isoparametric functions on Finsler space forms $*{}^{*}\,$

### v2

@@ 4.1 The number of distinct principal curvatures
Theorem 1.1 and Theorem 3.7 are not necessarily true for local transnormal functions. Obviously, the distance function to a hypersurface is a local transnormal function. But it is not necessarily isoparametric. Even for local isoparametric functions, The conclusion of Theorem 3.7 is not necessarily true.
@@
[16] Let $(\mathbb{R}^{3},F)$ be a 3-dimensional conic Minkowski-$(\alpha,\beta)$ space and the dual metric of $F$ be $F^{*}=\alpha^{*}\varphi(\frac{\beta^{*}}{\alpha^{*}})$, where
@@
where $\beta^{*}=(0,0,b)$ and $a,b$ are two positive constants. Then in $(\mathbb{R}^{3},F)$, helicoid $\mathbf{r}=(u\cos v,u\sin v,av)(0<u<1)$ is a local minimal isoparametric hypersurface with constant principal curvatures ±$\pm 1$.
In fact, we can construct a regular Finsler metric $\tilde{F}$ on $\mathbb{R}^{n}$ by the conic Finsler metric constructed in Lemma 4.1. Set $x_{0}=\frac{b}{(\varphi-s\varphi^{\prime})\sqrt{b^{2}-s^{2}}}$ and $x_{1}=x_{0}-\frac{1}{2a(\phi-s\phi^{\prime})\sqrt{b^{2}-s^{2}}}(s(\phi-s\phi^{\prime})+b^{2}\phi^{\prime})>x_{0}$. Denote
$L:\{\xi=(\xi_{1},0,\xi_{3})|F^{*}(\xi)=1,x_{0}<\xi<x_{1}\}.$
@@
$U=\{\xi\in\mathbb{R}^{3}~{}\big{|}~{}\xi_{1}^{2}+\xi_{2}^{2}>4a^{2}\xi_{3}^{2},\xi_{3}^{2}>0\}.$
Since $\frac{d\xi_{3}}{d\xi_{1}}\big{|}_{x_{0}}>0$, we can obtain a strictly convex smooth and closed surface $\tilde{\Sigma}$ including the origin in $\mathbb{R}^{3}$ by extending $\Sigma$.
We can define a regular Minkowski metric $\overline{F}^{*}$ satisfying $\overline{F}^{*}\mid_{\Sigma}=F^{*}$ on $\mathbb{R}^{3}$. Let $\overline{F}$ be the dual metric of $\overline{F}^{*}$. Obviously, $\overline{F}$ is still a regular Minkowski metric. In Lemma 4.1, the normal vector of the helicoid r is in $U$ if and only if $0<u<\frac{1}{2}$. Hence, $\textbf{r}=(u\cos v,u\sin v,av)$ ($0<u<\frac{1}{2}$) is a local isoparametric hypersurface in $(\mathbb{R}^{3},\overline{F})$ with constant principal curvature ±$\pm 1$.
Set $y_{1}=(y^{1},y^{2},y^{3})$, $y_{2}=(y^{4},\cdots,y^{n})$ and $y=(y_{1},y_{2})$. Then on $\mathbb{R}^{3}\times\mathbb{R}^{n-3}$ we define $\tilde{F}(y)=\sqrt{\overline{F}^{2}(y_{1})+|y_{2}|^{2}}$. The dual metric of $F$ is $\tilde{F}^{*}(\xi)=\sqrt{\overline{F}^{*2}(\xi_{1})+|\xi_{2}|^{2}}$, where $\xi_{1}=(\xi_{1},\xi_{2},\xi_{3})$, $\xi_{2}=(\xi_{4},\cdots,\xi_{n})$ and $\xi=(\xi_{1},\xi_{2})$. Let $\tilde{\textbf{r}}=(u\cos v,u\sin v,av,x_{2})$ be a hypersurface of $\mathbb{R}^{3}\times\mathbb{R}^{n-3}$, where $x_{2}=(x^{4},\cdots,x^{n})$. Then by a similar calculation in [16], the principal curvature is ±$\pm 1$ and 0. That is, when $0<u<\frac{1}{2}$, $\tilde{\textbf{r}}$ is a local isoparametric hypersurface in $(\mathbb{R}^{n},\tilde{F})$ with three principal curvatures, which is a level set of a local isoparametric function $f$. Hence, in the local case, we have
@@
In some Minkowski spaces, there exist local isoparametric hypersurfaces with three distinct principal curvatures at least.
@@ 4.2 Umbilic hypersurfaces
Let $(N,F)$ be a Finsler space with constant flag curvature. Then a totally umbilic hypersurface is local isoparametric.
@@
$^{i}\phi^{l}_{a}\phi^{p}_{b}(\frac{\partial\Gamma^{j}_{li}}{\partial x^{p}}(\textbf{n})+\frac{\partial\Gamma^{j}_{il}}{\partial y^{q}}(\textbf{n})\frac{\partial n^{q}}{\partial x^{p}})=n^{i}\phi^{l}_{a}\phi^{p}_{b}\frac{\delta\Gamma^{j}_{li}}{\delta x^{p}}(\textbf{n}).$ n ) .
Let $c$ be the constant flag curvature of $N$, then from [17],
@@
$D^{\textbf{n}}_{\partial u_{b}}D^{\textbf{n}}_{\partial u_{a}}\textbf{n}-D^{\textbf{n}}_{\partial u_{a}}D^{\textbf{n}}_{\partial u_{b}}\textbf{n}$ =$^{i}\phi^{l}_{a}\phi^{p}_{b}R^{j}_{ipl}(\textbf{n})=cn^{i}(\phi^{j}_{b}\phi^{l}_{a}g_{il}(\textbf{n})-\phi^{p}_{b}\phi^{l}_{a}g_{ip}(\textbf{n}))$ n ) )
@@
Hence, for any $a\neq b$, $\lambda$ is a constant. ∎
@@
From [7], we know
Lemma 5.1.
Let $(N,F)$ be an $n$-dimensional Finsler manifold with constant flag curvature. Then a transnormal function $f$ is isoparametric if and only if each regular level hypersurface of $f$ has constant principal curvatures.
Let $f$ be a transnormal function on a connected forward Finsler space form $(N(c),F)$ and $M$ be a totally umbilic level hypersurface of $f$. From Theorem 4.3, $M$ is local isoparametric. From (3.6), for any $t\in J$, $f^{-1}(t)$ has constant principal curvature. Combine Lemma 5.1, $f$ is isoparametric globally.
@@
Let $(N(c),F)$ be a forward Finsler space form, and $r(x)=d_{F}(p,x)$ be the distance function from $p$ to $x$. Set $f=\frac{1}{2}r^{2}$. Obviously, $f$ is a non-constant $C^{1}$ function on $\mathcal{D}_{p}$, which is smooth on $\mathcal{D}_{p}\setminus\{p\}$.
By a direct calculation, $F(\nabla f)=\sqrt{2f}$. From Laplace comparison theorem [18], we have $\hat{\Delta}r=(n-1)\frac{\mathfrak{s}^{\prime}_{c}(r)}{\mathfrak{s}_{c}(r)}$, where
@@
$\hat{\Delta}f=r\hat{\Delta}r+1=(n-1)\frac{\mathfrak{s}^{\prime}_{c}(r)}{\mathfrak{s}_{c}(r)}+1=\sqrt{2f}(n-1)\frac{\mathfrak{s}^{\prime}_{c}(\sqrt{2f})}{\mathfrak{s}_{c}(\sqrt{2f})}+1.$
Hence, $f$ satisfies (1.2) with $a(t)=2t,b(t)=\sqrt{2t}(n-1)\frac{\mathfrak{s}^{\prime}_{c}(\sqrt{2t})}{\mathfrak{s}_{c}(\sqrt{2t})}+1$. Therefore, $f$ is a local isoparametric function on $\mathcal{D}_{p}$.
@@
$\sum\limits_{a}\kappa_{a}^{2}=\left\{\begin{array}[]{lcl}\frac{1}{\sqrt{2}}(n-1)\frac{1}{t},&&{c=0,}\\ (n-1)c(\csc^{2}(\sqrt{2ct})-1),&&{c>0,}\\ -c(n-1)(\textmd{csch}^{2}(\sqrt{-2ct})+1),&&{c<0.}\end{array}\right.$ (6.14)
@@
$\kappa_{1}=\kappa_{2}=\ldots=\kappa_{n-1}=\left\{\begin{array}[]{lcl}-\frac{1}{\sqrt{2t}},&&{c=0,}\\ -\sqrt{c}\cot(\sqrt{2ct}),&&{c>0,}\\ -\sqrt{-c}\coth(\sqrt{-2ct}),&&{c<0,}\end{array}\right.$
@@
Let $(\mathbb{S}^{n}(c),F,d\mu)$ be a $n$-sphere equipped with a Finsler metric $F$ and the Busemann-Hausdorff volume form. If it has constant flag curvature and vanishing S-curvature, we call it a standard Finsler sphere[8]. We can obtain a global isoparametric function on $(\mathbb{S}^{n}(c),F,d\mu)$ with $g=1$. Set $f(x)=-\cos(\sqrt{c}r(x))$. From [8], we know that $f(x)$ is the first eigenvalue of $(\mathbb{S}^{n}(c),F,d\mu)$. Hence, from [19], $f\in C^{\infty}(\mathbb{S}^{n}_{f})$ and $f\in C^{1}(\mathbb{S}^{n})$. In addition, since $F(\nabla f)=\sqrt{c(1-f^{2})}$ and $\Delta f=-ncf$, then $f$ is isoparametric on $(\mathbb{S}^{n}(c),F,d\mu)$ with $a(t)=c(1-t^{2})$ and $b(t)=-nct$. From Lemma 6.1, we also have $\kappa_{1}=\kappa_{2}=\cdots=\kappa_{n-1}=\frac{\sqrt{c}t}{\sqrt{1-t^{2}}}$. The proof of Theorem 1.4 ends here.

### v3

@@ 4.1 The number of distinct principal curvatures
Theorem 1.1 and Theorem 3.9 are not necessarily true for locally transnormal functions. For example, the distance function to a hypersurface is a locally appropriate transnormal function for a reversible Finsler metric. But it is not necessarily isoparametric. Even for locally isoparametric functions, The conclusion of Theorem 3.9 is not necessarily true.
@@
[13] Let $(\mathbf{R}^{3},F)$ be a 3-dimensional conic Minkowski-$(\alpha,\beta)$ space and the dual metric of $F$ be $F^{*}=\alpha^{*}\varphi(\frac{\beta^{*}}{\alpha^{*}})$, where
@@
where $\beta^{*}=(0,0,b)$ and $a,b$ are two positive constants. Then in $(\mathbf{R}^{3},F)$, helicoid $\mathbf{r}=(u\cos v,u\sin v,av)(0<u<1)$ is a local minimal isoparametric hypersurface with constant principal curvatures $\pm 1$.
In fact, we can construct a regular Finsler metric $\tilde{F}$ on $\mathbf{R}^{n}$ by the conic Finsler metric constructed in Lemma 4.1. Set $x_{0}=\frac{b}{(\phi-s\phi^{\prime})\sqrt{b^{2}-s^{2}}}$ and $x_{1}=x_{0}-\frac{1}{2a(\phi-s\phi^{\prime})\sqrt{b^{2}-s^{2}}}(s(\phi-s\phi^{\prime})+b^{2}\phi^{\prime})>x_{0}$. Denote
$L:\{\xi=(\xi_{1},0,\xi_{3})|F^{*}(\xi)=1,x_{0}<\xi_{1}<x_{1}\}.$
@@
$U=\{\xi\in\mathbf{R}^{3}~\big{|}~\xi_{1}^{2}+\xi_{2}^{2}>4a^{2}\xi_{3}^{2},\xi_{3}^{2}>0\}.$
Since $\frac{d\xi_{3}}{d\xi_{1}}\big{|}_{x_{0}}>0$, we can obtain a strictly convex smooth and closed surface $\tilde{\Sigma}$ including the origin in $\mathbf{R}^{3}$ by extending $\Sigma$.
We can define a regular Minkowski metric $\overline{F}^{*}$ satisfying $\overline{F}^{*}\mid_{\Sigma}=F^{*}$ on $\mathbf{R}^{3}$. Let $\overline{F}$ be the dual metric of $\overline{F}^{*}$. Obviously, $\overline{F}$ is still a regular Minkowski metric. In Lemma 4.1, the normal vector of the helicoid r is in $U$ if and only if $0<u<\frac{1}{2}$. Hence, $\textbf{r}=(u\cos v,u\sin v,av)$ ($0<u<\frac{1}{2}$) is a locally isoparametric hypersurface in $(\mathbf{R}^{3},\overline{F})$ with constant principal curvature $\pm 1$.
[…]
@@
In some Minkowski spaces, there exist locally isoparametric hypersurfaces with three distinct principal curvatures at least.
From the analysis process of Theorem 4.2, we also know that
Remark 4.3.
There exist locally minimal isoparametric hypersurfaces in some Minkowski spaces, which are not Minkowski hyperplanes, Minkowski hyperspheres or Minkowski cylinders. Hence, locally isoparametric hypersurfaces may not be extended to global isoparametric hypersurfaces in Finsler manifolds, which is different from Riemannian case.
@@ 4.2 Umbilic hypersurfaces
Let $(N,F)$ be a Finsler space with constant flag curvature. Then a totally umbilic hypersurface is locally isoparametric.
@@
$n^{i}\phi^{l}_{a}\phi^{p}_{b}(\frac{\partial\Gamma^{j}_{li}}{\partial x^{p}}(\textbf{n})+\frac{\partial\Gamma^{j}_{il}}{\partial y^{q}}(\textbf{n})\frac{\partial n^{q}}{\partial x^{p}})=n^{i}\phi^{l}_{a}\phi^{p}_{b}\frac{\delta\Gamma^{j}_{li}}{\delta x^{p}}(\textbf{n}).$
Let $c$ be the constant flag curvature of $N$, then
@@
$D^{\textbf{n}}_{\partial u_{b}}D^{\textbf{n}}_{\partial u_{a}}\textbf{n}-D^{\textbf{n}}_{\partial u_{a}}D^{\textbf{n}}_{\partial u_{b}}\textbf{n}$ $=n^{i}\phi^{l}_{a}\phi^{p}_{b}R^{j}_{ipl}(\textbf{n})=cn^{i}(\phi^{j}_{b}\phi^{l}_{a}g_{il}(\textbf{n})-\phi^{p}_{b}\phi^{l}_{a}g_{ip}(\textbf{n}))$
@@
Hence, for any $a\neq b$, $\lambda$ is a constant. From Definition 2.4, $M$ is a locally isoparametric hypersurface. This completes the proof. ∎
@@
Let $f$ be a transnormal function on a connected forward Finsler space form $(N(c),F)$ and $M$ be a totally umbilic level hypersurface of $f$. Since $M$ is totally umbilic, from Theorem 4.4, the principal curvatures of $M$ are constants. The solution of (3.6) is determined by the principal curvatures of $M$ and the distance (3.1) between hypersurfaces. Hence, for any $t\in J$, $f^{-1}(t)$ has constant principal curvatures. Combine Definition 2.2, $f$ is a global isoparametric function on $(N(c),F)$.
Remark 5.1.
From the proof of Theorem 1.3, we know that if $f$ is a transnormal function on a forward Finsler space form, then the locally isoparametric hypersurfaces can be extended to the global isoparametric hypersurfaces.
@@
Let $(N(c),F)$ be a forward Finsler space form and $r(x)=d_{F}(p,x)$ be the distance function from $p$ to $x$. Set $f=\frac{1}{2}r^{2}$. Obviously, $f$ is a non-constant $C^{1}$ function on $\mathcal{D}_{p}$, which is smooth on $\mathcal{D}_{p}\setminus\{p\}$.
By a direct calculation, $F(\nabla f)=\sqrt{2f}$. From the proof of Theorem 4.1 in [16], we have $\hat{\Delta}r=(n-1)\frac{\mathfrak{s}^{\prime}_{c}(r)}{\mathfrak{s}_{c}(r)}$, where
@@
$\hat{\Delta}f=r\hat{\Delta}r+1=r(n-1)\frac{\mathfrak{s}^{\prime}_{c}(r)}{\mathfrak{s}_{c}(r)}+1=\sqrt{2f}(n-1)\frac{\mathfrak{s}^{\prime}_{c}(\sqrt{2f})}{\mathfrak{s}_{c}(\sqrt{2f})}+1.$
Hence, $f$ satisfies (1.2) with $a(t)=2t,b(t)=\sqrt{2t}(n-1)\frac{\mathfrak{s}^{\prime}_{c}(\sqrt{2t})}{\mathfrak{s}_{c}(\sqrt{2t})}+1$. Therefore, $f$ is a locally isoparametric function on $\mathcal{D}_{p}$.
@@
$\sum\limits_{a}\kappa_{a}^{2}=\left\{\begin{array}[]{lcl}\frac{1}{2t}(n-1),&&{c=0,}\\ (n-1)c(\csc^{2}(\sqrt{2ct})-1),&&{c>0,}\\ -c(n-1)(\textmd{csch}^{2}(\sqrt{-2ct})+1),&&{c<0.}\end{array}\right$. (6.14)
@@
$\kappa_{1}=\kappa_{2}=\ldots=\kappa_{n-1}=\left\{\begin{array}[]{lcl}-\frac{1}{\sqrt{2t}},&&{c=0,}\\ -\sqrt{c}\cot(\sqrt{2ct}),&&{c>0,}\\ -\sqrt{-c}\coth(\sqrt{-2ct}),&&{c<0.}\end{array}\right$.
@@
Let $\mathbb{S}^{n}$ be an $n$-sphere equipped with a Finsler metric $F$ and the Busemann-Hausdorff volume form $d\mu$. If it has constant flag curvature $c$, vanishing S-curvature and $Diam=\frac{\pi}{\sqrt{c}}$, we call it a standard Finsler sphere and denote it by $(\mathbb{S}^{n}(c),F,d\mu)$[19]. We can obtain a global isoparametric function on $(\mathbb{S}^{n}(c),F,d\mu)$ with $g=1$. Set $f(x)=-\cos(\sqrt{c}r(x))$. From [19], we know that $f(x)$ is the first eigenfunction of $(\mathbb{S}^{n}(c),F,d\mu)$. Hence, from [8], $f\in C^{\infty}(\mathbb{S}^{n}_{f})$ and $f\in C^{1}(\mathbb{S}^{n})$. In addition, since $F(\nabla f)=\sqrt{c(1-f^{2})}$ and $\Delta f=-ncf$, then $f$ is isoparametric on $(\mathbb{S}^{n}(c),F,d\mu)$ with $a(t)=c(1-t^{2})$ and b(t)=nctb(t)=-nct. From Lemma 6.1, we also have $\kappa_{1}=\kappa_{2}=\cdots=\kappa_{n-1}=\frac{\sqrt{c}t}{\sqrt{1-t^{2}}}$. That completes the proof of Theorem 1.4.